OpenStudy (anonymous):

A farmer has 7 and 3/4 rows of radishes in one field, 4 and 3/4 of radishes in another field, and 6 and 1/4 rows in a third field. How many rows of radishes does he have altogether?

OpenStudy (anonymous):

just add them all up.

OpenStudy (anonymous):

ok cause the answer key i have is showering 18 3/4 i wonder if they rounded up or something cause i know 7+4+6 =17 so that would of been 17 3/4

OpenStudy (anonymous):

You know that the whole numbers add up to 17. Now let's add the fraction: 3/4+3/4+1/4=7/4 (7/4 converted to mix number is 1 and 3/4) 17+1 3/4= 18 3/4

OpenStudy (anonymous):

ok